OpenStudy (hellomynameiskam):

Carla's plant was 10 2/3 inches tall. Yolanda's was 8 3/4 inches tall. How much taller was Carla's plant than Yolanda's? A. 1 11/12 in. B. 2 1/12 in. C. 2 5/10 in. D. 2 11/12

OpenStudy (hellomynameiskam):

i think its B.

OpenStudy (hellomynameiskam):

@KendrickLamar2014

OpenStudy (hellomynameiskam):

@AlexandervonHumboldt2

OpenStudy (hellomynameiskam):

@LegendarySadist

OpenStudy (anonymous):

The LCD of 3 and 4 is?

OpenStudy (hellomynameiskam):

3, 6, 9 , 12 4, 8 12

OpenStudy (anonymous):

So 12. Now, can you convert them to a denominator of 12?

OpenStudy (skullpatrol):

Are you sure option C is copied correctly. It should be 2 5/12, which is the answer, not 2 5/10 which would be 2 1/2 :-)

OpenStudy (hellomynameiskam):

C was a typo.. sorry @skullpatrol and yes

OpenStudy (anonymous):

Uhh, I don't think skullpatrol has that correct

OpenStudy (hellomynameiskam):

for C i meant to put 2 5/12 not 2 5/10

OpenStudy (anonymous):

Ok, but I was referring to skullpatrolling's implication that 2 5/12 is the answer. That isn't right.

OpenStudy (skullpatrol):

Convert the mixed numbers into improper fractions and use the common denominator of 12, then subtract.

OpenStudy (skullpatrol):

10 2/3 = 32/3 = 128/12 8 3/4 = 35/4 = 105/12 128/12 – 105/12 = 23/12 = 1 11/12 in.